# Single-qubit gates with errors at the $10^{-7}$ level

### v1

@@ Abstract
We report the achievement of single-qubit gates with sub-part-per-million error rates, in a trapped-ion 43Ca+ hyperfine clock qubit. We explore the speed/fidelity trade-off for gate times $4.4\leq t_{g}\leq 35~{}\upmu$s, and benchmark a minimum error of $1.5(4)\times 10^{-7}$. Gate calibration errors are suppressed to $<10^{-8}$, leaving qubit decoherence ($T_{2}\approx 70$ s), leakage and measurement as the dominant error contributions. The ion is held above a microfabricated surface-electrode trap which incorporates a chip-integrated microwave resonator for electronic qubit control; the trap is operated at room temperature without magnetic shielding.
@@
Electronically-controlled trapped ions have been used to perform the highest fidelity single- and two-qubit operations of any platform, with gate errors of $1\times 10^{-6}$ and $3\times 10^{-4}$ respectively [16, 3]. Even though single-qubit gates typically have higher fidelities than their two-qubit counterparts, quantum algorithms often require many more single-qubit than multi-qubit operations, and are therefore more sensitive to single-qubit gate errors [21]. High-fidelity single-qubit operations also have many other uses both in quantum information and further afield: for protecting “idle” qubits through dynamical decoupling [22]; in composite pulse sequences for error compensation [23] or addressing individual qubits [4]; and in quantum sensing applications [24, 25].
@@
Figure 1: Single-qubit gates: state-of-the-art. (a) State-of-the-art single-qubit gates across different quantum computing platforms. From left to right: electronically driven trapped ions [16], NV centres [26], laser-driven trapped ions [27], neutral atoms [28, 29], superconducting circuits [30] and semiconductor spin qubits [31]. (b) Single-qubit gate randomised benchmarking (RB) measurement carried out in this work. The average gate duration is $13\ \upmu$s; the fit yields an average Clifford gate error of $1.5(4)\times 10^{-7}$ and an average state-preparation and measurement (SPAM) error of $1.1(5)\times 10^{-3}$, which is consistent with independent SPAM measurements.
@@ B. Leakage, bit-flip and
Figure 5: Measurements of incoherent errors. (a) Decoherence errors are measured in a memory benchmarking experiment [38, 39] (orange circles), in which a delay $t$ is inserted between the $\pi/2$ pulses of a gate benchmarking sequence. A linear fit to the data (red) yields a decoherence time constant $T_{2}^{**}=69(7)$ s. This is in good agreement with the error measured at much longer timescales by a spin-echo Ramsey measurement [40, 41, 42, 20] (blue triangles). (b) The combined impact of leakage, bit-flip and time-dependent measurement errors is estimated using randomised benchmarking (RB) measurements where we replace all $\pi/2$ pulses with a delay of the same duration. Such measurements are collated here for delays corresponding to various gate times, and a linear fit is used to model this error.
Finally, we will discuss gate errors that are unrelated to MW pulse calibrations. From Table 1, we see that the dominant source of error is decoherence. This is measured in two ways: by memory benchmarking at short timescales [38, 39] and by spin-echo Ramsey measurement at long timescales [40, 41, 42, 20], with both sets of results shown in Fig. 5(a). During memory benchmarking, the delay between $\pi/2$ pulses in a RB sequence is increased and the subsequent increase in gate error provides a measure of a decoherence time constant, equal to $69(7)$ s, that we denote by $T_{2}^{**}$. The absence of the MW-induced ac Zeeman shift during the delay is accounted for by adjusting the phase of the subsequent $\pi/2$ pulses in the sequence. Errors caused by leakage, bit-flip and measurement are subtracted to isolate the effect of decoherence, using the independent measurement shown in Fig. 5(b). The memory benchmarking and Ramsey measurements are consistent, suggesting that the dominant decoherence mechanism occurs on time-scales similar to, or shorter than, the inter-pulse delays used in memory benchmarking ($\leq 150\ \upmu$s).
@@
In conclusion, we have achieved a new state-of-the-art single-qubit gate error of $1.5(4)\times 10^{-7}$ and presented a breakdown of all known sources of error, which agrees well with the data. For applications such as protecting memory qubits by dynamical decoupling, the relevant error would be the error per physical $\pi/2$ pulse, which is $0.7(2)\times 10^{-7}$. Our fastest Clifford gate ($4.4~{}\upmu$s) has an error of $2.9(5)\times 10^{-7}$; this represents a 20×20$\times$20 × speed-up and 3×3$\times$3 × lower error than the previous lowest-error gate [16]. Over the range of gate times studied, we determine the dominant contribution to the error to be decoherence, closely followed by leakage, bit-flip and time-dependent measurement errors. The gate error could therefore potentially be reduced further by improving the qubit coherence or by eliminating radiation leaking into the trap. Alternatively, the gate error could be reduced using state leakage detection techniques [45, 46, 47]. Calibrating the readout for different experimental sequence durations would allow us to separate the time-dependent measurement error from the gate error more unambiguously. Amplitude and frequency noise of the MW pulses contribute an insignificant amount to the gate error, demonstrating the robustness of electronic control.
Acknowledgments The authors thank A. Kwiatkowski, D.H. Slichter and R. Srinivas for insightful discussions. The trap was designed by T.P. Harty, with contributions from D.T.C. Allcock and D.M.L. This work was supported by the U.S. Army Research Office (ref. W911NF-18-1-0340) and the U.K. EPSRC Quantum Computing and Simulation Hub (ref. EP/T001062/1). M.C.S. acknowledges support from Balliol College, Oxford. A.D.L. acknowledges support from Oxford Ionics Ltd. K.M. acknowledges support from the Japan Science and Technology Agency (JST) ASPIRE Japan (ref. JPMJAP2319), and the JST Moonshot Research and Development program (ref. JPMJMS2063).
@@
In this section, we give further details on the randomised benchmarking (RB) results presented in Figs. 1,2. The ion is pseudo-randomly prepared in either state |$\ket{0}$ or |$\ket{1}$ before being subject to a pseudo-random series of Clifford gates. We then pseudo-randomly shelve either |$\ket{0}$ or |$\ket{1}$ into the “dark” $3^{2}\text{D}_{5/2}$ manifold and measure the probability of the ion fluorescing. The survival probability is defined as the probability of measuring the expected state at the end of the sequence. For each of the 5 different sequence lengths used in our measurements, we generate 30 pseudo-random gate sequences, and each is run 100 times. We refer to each of these 100 instances as a single “shot”. The outcomes from each of these measurements are shown in Fig. S2(a) for the optimal gate duration of 13 $\upmu$s. The complete dataset was taken over the course of 6 days, with data taken each day for all 10 different gate durations (results for all gate durations measured are shown in Fig. 2 and Fig. S2(b)). RB measurements were repeated but with all microwave pulses replaced by delays of the same duration to measure the leakage, bit-flip and time-dependent measurement error.
Figure S2: (a) Raw data underlying the RB measurement result shown in Fig. 1. For each of the 5 sequence lengths used in our measurements, we generate 30 pseudo-random gate sequences and run each one 100 times. The histogram shows the percentage of times the sequence failed for all 30 gate sequences. Results are shown for gate RB measurements (blue) and also for “idle RB” measurements (orange), where all gate pulses are replaced by delays of the same duration. (b) Gate and idle RB measurement results for all gate durations. Gate RB errors resulting from these measurements are plotted in Fig. 2, and idle RB errors are shown in Fig. 5(b).
@@
where $l$ is the sequence length, $\epsilon$ is the depolarisation rate, and $A$ is a fit parameter which accounts for SPAM errors [34]. The error per Clifford gate is then given by $\epsilon$. Given the binary nature of the measurement outcomes (fluorescence or no fluorescence), we assume that the data follows a binomial distribution. For each randomised sequence, we calculate the likelihood of obtaining the observed measurement outcome, and MLE is used to find the best-fitting parameters that maximise this likelihood. To estimate the uncertainties in the fitted parameters, we employ parametric bootstrapping. Using the fitted parameters from the MLE, we generate simulated datasets that follow the same binomial distribution. The fitting procedure is repeated on these simulated datasets, yielding a distribution of the fit parameters. The standard deviation of this distribution is then used to determine the uncertainties on our final reported values.
@@
The data acquired through memory benchmarking, shown in Fig. 5(a), is fitted following
@@
The presence of these states can affect the gate error in two ways. Firstly, population can remain within these states at the end of a pulse, causing leakage out of the qubit subspace. The impact of this is determined by the pulse ramp time (40 ns in our case) relative to the effective Rabi period for these spectator transitions ($\approx 1/\Delta_{s}\approx 10$ ns). Secondly, the portion of the population residing in these spectator states is subject to increased decoherence, resulting from the first-order magnetic field dependence of the spectator transitions. We measure a decoherence time $T_{2,S}^{**}\sim 40$ ms for these transitions using memory benchmarking, which is three orders of magnitude less than the qubit coherence time. However, this reduced coherence affects only a small proportion of the population residing in each state $\sim\Omega_{S}^{2}/(\Delta_{S}^{2}+\Omega_{S}^{2})\lesssim 3\times 10^{-6}$.
@@
|$\begin{split}\ket{\psi}&=\exp\left(-\frac{\text{i}}{2}\Omega_{q}A_{0}\hat{\sigma}_{x}+\sum_{k=0}^{\infty}-\frac{\text{i}}{2}\Omega_{k}A_{k}\hat{\sigma}_{x}\right)\ket{0}\ \text{and}\ P_{\ket{0}}=\cos^{2}\left(\sum_{k=0}^{\infty}\frac{\Omega_{k}A_{k}}{2}\right)\ ,\\ A_{k}&=\frac{1}{\left(k+1\right)}\left(\frac{2\pi N}{\Omega_{q}2^{M}}\right)^{k+1}\sum_{m=0}^{M}W(m)\left(\left(m+1\right)^{k+1}-m^{k+1}\right)\ ,\end{split}$ (S21)
@@
S5 Leakage, bit-flip and time-dependent measurement errors
@@
Figure S4: Diagnostic measurements of measurement, leakage and bit-flip errors. In all measurements presented, we prepare the ion in a qubit state (|$\ket{0}$ or |$\ket{1}$), wait for a varying delay, and shelve a chosen set of states which will be rendered dark in state-dependent fluorescence readout. The shelving scheme varies across measurements as follows: in (a) we measure the probability of the ion fluorescing after a delay, without shelving any states; in (b) we shelve the initially prepared state; in (c) we shelve the qubit state that was not prepared; and in (d) we shelve both qubit states. For each type of measurement, we measure after preparing |$\ket{0}$ (blue) or after preparing |$\ket{1}$ (red). The measurement, leakage and bit-flip error contributions for each measurement are shown in the formulas on the right. For (b), we also show the same measurement, starting in |$\ket{0}$, before certain improvements were made to reduce leakage errors (grey).
Figure S5: Improvements in bit-flip errors. By combining the measurements of Fig S4, we can extract the bit-flip error probability $P_{0\leftrightarrow 1}$ as a function of delay time as described in Sec. S5. Here we show the resulting error before and after the addition of attenuation after the microwave amplifier (see Fig. S1 for the microwave drive chain).

### v2

@@ Abstract
We report the achievement of single-qubit gates with sub-part-per-million error rates, in a trapped-ion 43Ca+ hyperfine clock qubit. We explore the speed/fidelity trade-off for gate times $4.4\leq t_{g}\leq 35~{}\upmu$s, and benchmark a minimum error per Clifford gate of $1.5(4)\times 10^{-7}$. Calibration errors are suppressed to $<10^{-8}$, leaving qubit decoherence ($T_{2}\approx 70$ s), leakage, and measurement as the dominant error contributions. The ion is held above a microfabricated surface-electrode trap which incorporates a chip-integrated microwave resonator for electronic qubit control; the trap is operated at room temperature without magnetic shielding.
@@
Electronically-controlled trapped ions have been used to perform the highest fidelity single- and two-qubit operations of any platform, with gate errors of $1.0(3)\times 10^{-6}$ and $3(1)\times 10^{-4}$ respectively [16, 3]. Even though single-qubit gates typically have higher fidelities than their two-qubit counterparts, quantum algorithms often require many more single-qubit than multi-qubit operations, and are therefore more sensitive to single-qubit gate errors [21]. High-fidelity single-qubit operations also have many other uses both in quantum information and further afield: for protecting “idle” qubits through dynamical decoupling [22]; in composite pulse sequences for error compensation [23] or addressing individual qubits [4, 3]; and in quantum sensing applications [24, 25].
@@
Figure 1: Single-qubit gates: state-of-the-art. (a) State-of-the-art single-qubit gates across different quantum computing platforms. From left to right: electronically driven trapped ions [16], NV centres [26], laser-driven trapped ions [27], superconducting circuits [28], neutral atoms [29, 30], and semiconductor spin qubits [31]. (b) Single-qubit gate randomised benchmarking (RB) measurement carried out in this work. The average gate duration is $13\ \upmu$s; the fit yields an average Clifford gate error of $1.5(4)\times 10^{-7}$ and an average state-preparation and measurement (SPAM) error of $1.1(5)\times 10^{-3}$, which is consistent with independent SPAM measurements.
@@ B. Leakage, bit-flip and
Figure 5: Measurements of incoherent errors. (a) Decoherence errors are measured in an interleaved randomised memory benchmarking experiment [38, 39] (orange circles), in which a delay $t$ is inserted between the $\pi/2$ pulses of a gate benchmarking sequence. A linear fit to the data (red) yields a decoherence time constant $T_{2}^{**}=69(7)$ s. This is in good agreement with the error measured at much longer timescales by a spin-echo Ramsey measurement [40, 41, 42, 20] (blue triangles). (b) The combined impact of leakage, bit-flip and time-dependent measurement errors is estimated using randomised benchmarking (RB) measurements where we replace all $\pi/2$ pulses with a delay of the same duration. Such measurements are collated here for delays corresponding to various gate times, and a linear fit is used to model this error.
Finally, we will discuss gate errors that are unrelated to MW pulse calibrations. From Table 1, we see that the dominant source of error is decoherence. This is measured in two ways: by interleaved randomised memory benchmarking (IRMB) at short timescales [38, 39] and by spin-echo Ramsey measurement at long timescales [40, 41, 42, 20], with both sets of results shown in Fig. 5(a). During IRMB, the delay between $\pi/2$ pulses in a RB sequence is increased and the subsequent increase in gate error provides a measure of a decoherence time constant, equal to $69(7)$ s, that we denote by $T_{2}^{**}$. The absence of the MW-induced ac Zeeman shift during the delay is accounted for by adjusting the phase of the subsequent $\pi/2$ pulses in the sequence. Errors caused by leakage, bit-flip and measurement are subtracted to isolate the effect of decoherence, using the independent measurement shown in Fig. 5(b). We determine the dominant source of decoherence to be phase noise in the MW drive chain [43, 44] (further analysis can be found in [32, §S5]).
@@
In conclusion, we have achieved a new state-of-the-art single-qubit gate error of $1.5(4)\times 10^{-7}$ and presented a breakdown of all known sources of error, which agrees well with the data over a wide range of gate speeds. Our fastest Clifford gate ($4.4~{}\upmu$s) has an error of $2.9(5)\times 10^{-7}$; this represents a 20×20$\times$20 × speed-up relative to the previous lowest-error gate [16]. Compared with our previous work [16], we have reduced the error by nearly an order of magnitude by better control of microwave amplitude and detuning with automated calibration procedures, and by reduced excitation of spectator transitions using pulse shaping and larger Zeeman splittings. Over the range of gate times studied, we determine the dominant contribution to the error to be decoherence, closely followed by leakage, bit-flip and time-dependent measurement errors. The gate error could therefore be reduced further by improving the qubit coherence or by eliminating leakage out of the qubit states due to unwanted sources of radiation. Alternatively, the gate error could be reduced using state leakage detection techniques [47, 48, 49]. Calibrating the readout for different experimental sequence durations would allow us to separate the time-dependent measurement error from the gate error more unambiguously. Amplitude and frequency noise of the MW pulses contribute an insignificant amount to the gate error, demonstrating the robustness of electronic control. With hardware improvements, one could implement longer gate sequences, enabling experiments to probe the non-Markovian nature of the different sources of error [50, 51, 52, 53]. Note that these types of errors could cause the RB measurements used here (and in many other similar experiments) to underestimate the gate error [54, 55]. For addressed single-qubit gates, multiple techniques exist to null the microwave field in non-addressed zones [56, 57, 3], and even modest levels of cancellation can result in very low crosstalk through the use of composite pulses [58, 3]. Alternatively, this gate mechanism can drive different single qubit gates on $n$ ions within the same zone or potential well, using a sequence of $\sim 3n/2$ pulses [4]. For composite pulse sequences, or applications such as protecting memory qubits by dynamical decoupling, the relevant error would be the error per physical $\pi/2$ pulse, which is $0.7(2)\times 10^{-7}$.
Acknowledgments The authors wish to thank A. Kwiatkowski, D.H. Slichter and R. Srinivas for insightful discussions; and J. Grover and H. Stemp for comments on the manuscript. The trap was designed by T.P. Harty, with contributions from D.T.C. Allcock and D.M.L. This work was supported by the U.S. Army Research Office (ref. W911NF-18-1-0340) and the U.K. EPSRC Quantum Computing and Simulation Hub (ref. EP/T001062/1). M.C.S. acknowledges support from Balliol College, Oxford. A.D.L. acknowledges support from Oxford Ionics Ltd. K.M. acknowledges support from the Japan Science and Technology Agency (JST) ASPIRE Japan (ref. JPMJAP2319), and the JST Moonshot Research and Development program (ref. JPMJMS2063).
@@
In this section, we give further details on the randomised benchmarking (RB) results presented in Figs. 1,2. The ion is pseudo-randomly prepared in either state |$\ket{0}$ or |$\ket{1}$ before being subject to a pseudo-random series of Clifford gates. We then pseudo-randomly shelve either |$\ket{0}$ or |$\ket{1}$ into the “dark” $3^{2}\text{D}_{5/2}$ manifold and measure the probability of the ion fluorescing. The survival probability ($1-$ error probability) is defined as the probability of measuring the expected state at the end of the sequence. For each of the 5 different sequence lengths used in our measurements, we generate 30 pseudo-random gate sequences (generated using [59], list of sequences is available on request), and each is run 100 times. We refer to each of these 100 instances as a single “shot”. The outcomes from each of these measurements are shown in Fig. S2(a) for the optimal gate duration of 13 $\upmu$s. The complete dataset was taken over the course of 6 days, with data taken each day for all 10 different gate durations (results for all gate durations measured are shown in Fig. 2 and Fig. S2(b)). RB measurements were repeated but with all microwave pulses replaced by delays of the same duration to measure the leakage, bit-flip and time-dependent measurement error.
Figure S2: (a) Raw data underlying the RB measurement result shown in Fig. 1. For each of the 5 sequence lengths used in our measurements, we generate 30 pseudo-random gate sequences and repeat each one 100 times. The histogram shows the number of times an error was measured for all 30 gate sequences. Results are shown for gate RB measurements (blue) and also for “idle RB” measurements (orange), where all gate pulses are replaced by delays of the same duration. (b) Gate and idle RB measurement results for all gate durations. Gate RB errors resulting from these measurements are plotted in Fig. 2, and idle RB errors are shown in Fig. 5(b).
@@
where $l$ is the sequence length, $\epsilon$ is the depolarisation rate, and $A$ is a fit parameter which accounts for SPAM errors [34]. The error per Clifford gate is then given by $\epsilon$.
This model assumes that errors are Markovian which may not always be true. A more rigorous analysis would involve fitting the data to a more general model, such as those presented in Refs. [50, 51]. Since, in our measurements, the total error we accumulate at the maximum applied sequence length is only $4.7\times 10^{-3}$ (the maximum sequence length is limited by our hardware), we do not observe a significant deviation from the simple exponential model and so cannot investigate the presence of non-Markovian errors using our RB data.
Given the binary nature of the measurement outcomes (fluorescence or no fluorescence), we assume that the data follows a binomial distribution. For each randomised sequence, we calculate the likelihood of obtaining the observed measurement outcome, and MLE is used to find the best-fitting parameters that maximise this likelihood. To estimate the uncertainties in the fitted parameters, we employ parametric bootstrapping. Using the fitted parameters from the MLE, we generate simulated datasets that follow the same binomial distribution. The fitting procedure is repeated on these simulated datasets, yielding a distribution of the fit parameters. The standard deviation of this distribution is then used to determine the uncertainties on our final reported values.
@@
The data acquired through interleaved randomised memory benchmarking, shown in Fig. 5(a), is fitted following
@@
The presence of these states can affect the gate error in two ways. Firstly, population can remain within these states at the end of a pulse, causing leakage out of the qubit subspace. The impact of this is determined by the pulse ramp time (40 ns in our case) relative to the effective Rabi period for these spectator transitions ($\approx 1/\Delta_{s}\approx 10$ ns). Secondly, the portion of the population residing in these spectator states is subject to increased decoherence, resulting from the first-order magnetic field dependence of the spectator transitions. We measure a decoherence time $T_{2,S}^{**}\sim 40$ ms for these transitions using interleaved randomised memory benchmarking, which is three orders of magnitude less than the qubit coherence time. However, this reduced coherence affects only a small proportion of the population residing in each state $\sim\Omega_{S}^{2}/(\Delta_{S}^{2}+\Omega_{S}^{2})\lesssim 3\times 10^{-6}$.
@@
|$\begin{split}\ket{\psi}&=\exp\left(-\frac{\text{i}}{2}\Omega_{q}A_{0}\hat{\sigma}_{x}+\sum_{k=0}^{\infty}-\frac{\text{i}}{2}\Omega_{k}A_{k}\hat{\sigma}_{x}\right)\ket{0}\ \text{and}\ P_{\ket{0}}=\cos^{2}\left(\sum_{k=0}^{\infty}\frac{\Omega_{k}A_{k}}{2}\right)\ ,\\ A_{k}&=\frac{1}{\left(k+1\right)}\left(\frac{2\pi N}{\Omega_{q}2^{M}}\right)^{k+1}\quad\sum_{m=0}^{2^{M}-1}W(m)\left(\left(m+1\right)^{k+1}-m^{k+1}\right)\ ,\end{split}$ (S21)
@@
S5 Sources of Decoherence
In this section, we discuss the different sources of decoherence present in our experimental setup and compare them to the measurements in Fig. 5(a).
S5.1 Phase noise in the microwave drive chain
Figure S4: Phase noise in the microwave drive chain. (a) A simplified schematic of the microwave drive chain used to generate pulses at the qubit frequency. A more detailed schematic can be found in Fig. S1(b). (b) The single-sideband (SSB) phase noise measured at different points along the microwave drive chain, indicated by stars in (a). At low frequencies ($<1$ Hz) the phase noise is dominated by the local oscillator and at high frequencies ($>10$ kHz) the phase noise of the amplifier determines the noise floor, which is well above the level set by thermal noise (dashed line). Filter-transfer functions describing (c) interleaved randomised memory benchmarking (IRMB) and (d) spin-echo Ramsey measurements, were used in combination with the SSB phase noise measurements to determine the expected error during these measurements (red triangles). Comparing to experimental results (blue circles), shown in Fig. 5(a), indicates that phase noise plays a significant role in our observed decoherence. The phase noise measurements lead to excellent agreement in the low inter-pulse delay regime of IRMB, closest to the RB measurement. Other effects could lead to inflated errors at longer delays such as a slight miscalibration of the bare qubit frequency by $\sim 2.5$ Hz; accounting for this detuning leads to better agreement with the IRMB measurement at all delays (black crosses).
Previous studies have shown that phase noise is the dominant source of decoherence in many trapped-ion experiments [39, 43, 20, 65]. To determine the phase noise in our microwave drive chain, we measured the single-sideband (SSB) phase noise at different points along the chain using a phase noise analyser (Rohde &\&& Schwarz FSWP), as shown in Fig. S4. At low frequencies ($<1$ Hz) the phase noise is dominated by noise from the local oscillator. We use a multiplied crystal oscillator, which converts a $10$ MHz Rb clock signal to $3.00$ GHz. For an ideal oscillator, this process would amplify the SSB phase noise in the clock by $20\log_{10}\left(\omega_{\text{LO}}/\omega_{\text{clock}}\right)=50~{}$dB [66], which is consistent with the measured local oscillator phase noise. The low-frequency noise could therefore be reduced by improving clock phase noise, and/or reducing the noise amplification by using type of oscillator, for example a cryogenic sapphire oscillator [67, 68].
At high frequencies ($>10$ kHz) the phase noise of the amplifier dominates the noise. We determine the white noise floor set by the amplifier to be well above the level set by room-temperature thermal noise (electronic noise generated by the thermal movement of charge carriers), indicated by the dashed line in Fig. S4(b). The thermal noise floor of the SSB phase noise (in units of dBc/Hz) is calculated using
$\mathcal{L}_{\text{min}}=30+10\text{log}\left(k_{\text{B}}T\right)-P_{\text{carrier}}$ (S24)
where $k_{\text{B}}$ is Boltzmann’s constant, $T$ is the temperature in K, and $P_{\text{carrier}}$ is the power of the carrier frequency in units of dBm [43, 69].
The influence of phase noise on coherence measurements can be determined using filter transfer functions [43, 44, 69]. The measured SSB phase noise must first be transformed into a (unilateral) phase power spectral density
$\mathcal{S}_{\phi}\left(\omega\right)=2\times 10^{\mathcal{L}\left(\omega\right)/10}$ (S25)
where $\mathcal{L}\left(\omega\right)$ is the SSB phase noise [70, 69]. The fidelity of a sequence of qubit operations as a function of time is then determined by
$\mathcal{F}\left(\tau\right)=\frac{1}{2}\left(1+e^{-\chi\left(\tau\right)}\right)$ (S26)
where $\chi\left(\tau\right)$ is the overlap integral of the phase power spectral density with a filter transfer function describing the sequence of qubit operations
[…]
We used an existing Python package to generate filter transfer functions [71] describing the decoherence measurements presented in Fig. 5(a). A description of how to derive filter transfer functions for sequences of unitary qubit operations can be found in Refs. [71, 72].
To simulate IRMB measurements, we generated filter transfer functions to model different random sequences of up to $10,000$ Clifford gates. We then computed the expected error for each sequence using Eq. S27, and fitted the results using Eq. S1 to determine the IRMB error shown in Fig. S4(c). For spin-echo Ramsey measurements, we generated filter transfer functions for a range of delays and computed the expected error using Eq. S27, shown in Fig. S4(d). Filter transfer functions for spin-echo experiments can also be written analytically [43, 44], which we used to verify our model. Comparing the resulting “simulated” data indicates that phase noise in the microwave drive chain is a significant source of decoherence in our system.
S5.2 Magnetic field fluctuations
Another possible source of decoherence is fluctuations in the static magnetic field. To assess their impact, we first measured the frequency of the qubit transition as a function of magnetic field, confirming we are operating at a field where the qubit has a minimal first-order sensitivity to “fast” magnetic field fluctuations. This also allows us to determine the change in sensitivity in the event of “slow” drift in the magnetic field mean value. We then measured the decoherence time $T_{2}^{*}$ of a first-order magnetic field sensitive “stretch” transition (|$\ket{F=4,M=4}\leftrightarrow\ket{F=3,M=3}$), using a spin-echo Ramsey measurement, to be $1.9(2)$ ms, revealing “fast” fluctuations on the order of $\sim 10$ nT. Between each of our RB measurements, we also calibrate the static magnetic field by measuring the change in the frequency of the stretch transition, revealing “slow” changes in the magnetic field of $\sim 80$ nT. These “fast” and “slow” changes in the magnetic field, combined with measurements of the qubit sensitivity, provide an estimate of the decoherence time as $\sim 3700$ s, making these effects negligible relative to the measured decoherence and the decoherence induced by microwave phase noise.
S6 Leakage, bit-flip and time-dependent measurement errors
@@
Figure S5: Diagnostic measurements of measurement, leakage and bit-flip errors. In all measurements presented, we prepare the ion in a qubit state (|$\ket{0}$ or |$\ket{1}$), wait for a varying delay, and shelve a chosen set of states which will be rendered dark in state-dependent fluorescence readout. The shelving scheme varies across measurements as follows: in (a) we measure the probability of the ion fluorescing after a delay, without shelving any states; in (b) we shelve the initially prepared state; in (c) we shelve the qubit state that was not prepared; and in (d) we shelve both qubit states. For each type of measurement, we measure after preparing |$\ket{0}$ (blue) or after preparing |$\ket{1}$ (red). The measurement, leakage and bit-flip error contributions for each measurement are shown in the formulas on the right. For (b), we also show the same measurement, starting in |$\ket{0}$, before certain improvements were made to reduce leakage errors (grey). These measurements were combined to estimate the probability of a bit-flip error occurring, see Fig. S6.
Figure S6: Improvements in bit-flip errors. By combining the measurements of Fig S5, we can extract the bit-flip error probability $P_{0\leftrightarrow 1}$ as a function of delay time as described in Sec. S6. Here we show the resulting error when the attenuation was placed before or after the microwave amplifier (see Fig. S1 for the microwave drive chain).